Solve each system using the specified method. Substitution
4x-3y=15
x-2y=0

Found 2 solutions by checkley77, solver91311:
Answer by checkley77(12569)

4x-3y=15
x-2y=0 x=2y now replace x with (2y) in the first equation & solve for y.
4(2y)-3y=15
8y-3y=15
5y=15
y=15/5
y=3 ans.
x=2*3
x=6 ans.
Proof:
4*6-3*3=15
24-9=15
15=15

Answer by solver91311(12126)






Solve equation 2. for x:



Substitute 2y for x in equation 1.:



Now solve for y. Once you have found the value for y, substitute that value into either of the original equations to solve for x. I suggest you use equation 2 for this step.

John